# Interpretable Multimodal Learning for Tumor Protein-Metal Binding: Progress, Challenges, and Perspectives

### v1

@@ Abstract
In cancer therapeutics, protein-metal binding mechanisms critically govern drug pharmacokinetics and targeting efficacy, thereby fundamentally shaping the rational design of anticancer metallodrugs. While conventional laboratory methods used to study such mechanisms are often costly, low throughput, and limited in capturing dynamic biological processes, machine learning (ML) has emerged as a promising alternative. Despite increasing efforts to develop protein–metal binding datasets and ML algorithms, the application of ML in tumor protein-metal binding remains limited. Key challenges include a shortage of high-quality, tumor-specific datasets, insufficient consideration of multiple data modalities, and the complexity of interpreting results due to the “black box” nature of complex ML models. This paper summarizes recent progress and ongoing challenges in using ML to predict tumor protein-metal binding, focusing on data, modeling, and interpretability. We present multimodal protein-metal binding datasets and outline strategies for acquiring, curating, and preprocessing them for training ML models. Moreover, we explore the complementary value provided by different data modalities and examine methods for their integration. We also review approaches for improving model interpretability to support more trustworthy decisions in cancer research. Finally, we offer our perspective on research opportunities and propose strategies to address the scarcity of tumor protein data and the limited number of predictive models for tumor protein–metal binding. We also highlight two promising directions for effective metal-based drug design: integrating protein–protein interaction data to provide structural insights into metal-binding events and predicting structural changes in tumor proteins after metal binding.
@@
  • Cross-referencing datasets can help address the scarcity of tumor protein metal-binding data.
@@ 1 Introduction
The design and characterization of metal-based compounds mainly rely on conventional wet-lab methods, such as mass spectrometry, X-ray crystallography and nuclear magnetic resonance [5, 6, 7, 8], as shown in Fig. 1A. These methods provide valuable insights into the properties of metal-based compounds and the structural basis of their interactions with tumor proteins. However, they are often expensive, time-consuming, and low throughput, and they struggle to capture dynamic biological processes. The emergence of machine learning (ML) presents new opportunities, as ML techniques can efficiently analyze large amounts of complex biological data and build predictive models for protein-metal binding, thereby accelerating metal-based drug discovery. Figure 1D shows three key ML application tasks in tumor protein-metal binding: tumor metalloprotein identification, metal type classification, and binding site prediction. Despite this potential, the application of ML remains limited due to challenges including the scarcity of high-quality datasets, insufficient integration of multimodal information, and lack of interpretability in existing models.
@@
Figure 1: Interpretable multimodal learning for tumor protein-metal binding. A) The process of transferring bioinformatics data to ML-oriented datasets. B) The data modalities, their combinations and the commonly-used multimodal learning methods. C) The key roles of model interpretability. D) The primary applications for cancer research. Created with BioRender.
@@
To summarize recent advances, highlight key challenges, and explore promising directions, this paper is organized around three key areas: data resources, multimodal learning approaches, and interpretability techniques. First, we reviewed existing data resources and observed that, while high-quality general protein-metal binding datasets are abundant, those specific to tumor proteins are scarce. Since tumor proteins are not structurally different from general proteins, the learning mechanisms in ML models remain applicable across both types. Therefore, we collected datasets on both general protein-metal binding and tumor proteins. We then proposed a strategy to construct tumor-specific datasets by integrating these complementary sources. For example, although the Protein Data Bank (PDB) [9] does not explicitly annotate tumor protein-metal binding interactions, it includes cancer-associated proteins with experimentally resolved metal-binding structures. Identifying and using these entries to extract key biological features can help mitigate current data scarcity.
@@
Next, different data modalities can provide complementary biological insights to enhance feature learning when effectively integrated. To capture this diversity, we identified four key data modalities relevant to tumor protein: sequence, structure, protein pocket, and text, as shown in Fig. 1B. For each modality, we introduced its unique characteristics and outlined an ML-oriented data processing workflow, including data acquisition, data preprocessing and feature extraction. This workflow can support researchers in locating, preparing, and adopting raw data for ML model development.
@@
Integrating these heterogeneous modalities into a model is nontrivial and can significantly influence model performance. To better understand the current landscape, we examined existing ML models developed for tumor protein-metal binding. To the best of our knowledge, only three such models have been developed to date - multichannel convolutional neural network (MCCNN) [10], MetalPrognosis [11], and MetalTrans [12] - but they all have limited success in multimodal integration. Therefore, we studied multimodal ML methods from broader bioinformatics applications, including 14 recent protein–metal binding models and four protein foundation models, and proposed strategies for adapting them to tumor-specific tasks. Building on these insights, we highlighted the important value of multimodality for tumor protein-metal binding, suggested promising combinations of data modalities, and provided an overview of general multimodal learning methods, as shown in Fig. 1B.
@@
Finally, in medical and biological applications, predictive accuracy alone is insufficient. ML models should also be interpretable and trustworthy to support scientific and clinical understanding. However, the “black box” nature of popular complex ML models limits transparency into their internal decision-making processes. In practice, interpretability is often restricted to indirect evaluations, such as performance comparisons with benchmark models, without clarifying how specific inputs contribute to predictions. This lack of transparency poses challenges for researchers, particularly those from clinical and biological backgrounds, in assessing model reliability and extracting clinically or biologically relevant insights. To address this issue, we examined interpretability in the context of tumor protein-metal binding from two perspectives: designing models with inherent interpretability and applying post-hoc analysis techniques.
@@
Building on these findings, we summarized key challenges and discussed the perspectives on potential solutions. Moreover, we identified two emerging directions that can advance tumor protein–metal binding prediction. The first involves integrating protein-protein interaction data to provide a structural context for metal-binding events. The second focuses on predicting the structural changes in tumor proteins after metal binding, which can inform the design of more effective metal-based therapeutics. To our knowledge, this is the first comprehensive study of interpretable multimodal ML for tumor protein-metal binding. Our work synthesized research advances, identified ongoing challenges, and offered a forward-looking perspective into future research opportunities. By integrating knowledge from protein structure datasets, multimodal ML, and interpretability analysis, we aim to support the development of biologically grounded, accurate, and clinically applicable ML models for tumor protein-metal binding prediction.
The main contributions of this paper can be summarized as follows:
@@
  • We presented and summarized potential data resources for tumor protein-metal binding, outlined a systematic data processing workflow, and proposed a strategy to construct tumor-specific datasets by integrating protein-metal binding and cancer-related datasets.
@@
  • We identified key protein data modalities, explored their combinations, and reviewed multimodal learning methods focusing on their applicability to tumor protein-metal binding. We also emphasized the importance of interpretability and summarized current approaches that support biological and clinical insight.
@@
  • We offered a forward-looking perspective by proposing two promising research directions, i.e., exploring protein-protein interaction and predicting tumor protein structural changes upon metal binding, to inform drug design and discovery.
@@
The remainder of this paper is organized as follows. Section 2 introduces the role of metal ions in tumor proteins. Section 3 outlines a workflow for processing multimodal data. Section 4 introduces strategies for combining data modalities and reviews existing multimodal learning methods. Section 5 introduces current efforts to improve model interpretability. Section 6 highlights two key considerations for advancing future research in tumor protein-metal binding. Finally, Section 7 presents our conclusions.
@@ 2 Biological significance of tumor protein-metal binding
These protein-metal interactions form a highly coordinated regulatory system that influences tumor progression, immune response, and cellular metabolism. The biological significance underscores the potential for developing targeted cancer therapies and motivates this work.
3 Tumor protein-metal binding data
Given the limited availability of tumor-specific protein-metal binding datasets, applying ML in this domain remains challenging. This section reviewed existing data resources, identified key data modalities, and proposed strategies to mitigate data scarcity. We introduced publicly available cancer-related and protein-metal binding datasets and discussed how integrating these resources can support the construction of tumor-specific datasets. Additionally, we examined four key data modalities in protein research and outlined a data processing workflow to prepare these modalities for ML model development.
3.1 Tumor protein and metal-binding datasets
Several publicly available datasets offer valuable information on cancer biology and protein-metal binding, enabling researchers to investigate the structural and functional roles of metal ions in both tumor and general proteins. As summarized in Table 1 and 2, cancer-related datasets mainly focus on genomic, proteomic, and clinical data relevant to tumor biology, while protein-metal binding datasets contain detailed structural and functional annotations for proteins that coordinate metal ions. Because tumor proteins are not structurally distinct from general proteins, we propose a strategy to integrate tumor protein annotations with metal-binding data, enabling researchers to cross-reference information across specialized datasets. By leveraging multiple data sources, researchers can identify proteins of interest and explore their metal coordination properties and relevance in tumor biology. For example, a protein can be identified through the Cancer Genome Atlas Program (TCGA) [22], and its metal-binding features, such as coordination details, binding motifs, and potential roles, can be explored using UniProt [23]. A case in point is the copper transport protein ATOX1, which exhibits altered RNA transcript levels in various cancers [24]. Researchers can find ATOX1 in TCGA and use its UniProt accession number (O00244) to access its metal-binding information. ATOX1 binds copper and functions as a chaperone, delivering copper to targets such as P-type ATPases. The dataset also links to studies on ATOX1’s roles in breast and liver cancer, while structural data is accessible via external resources like PDB.
This strategy highlights how integrating datasets such as TCGA and PDB enables researchers to investigate metal-binding proteins, even when explicit metal annotations are missing in tumor datasets, by leveraging cross-referenced structural and functional insights. Looking ahead, we envision a future where an integrated dataset combines gene expression, metal-binding, and structural data, streamlining protein-level analysis in cancer. Such a resource would allow researchers to readily query proteins like ATOX1 to assess tumor-specific expression, metal coordination, and structural context, ultimately supporting the discovery of new targets for metal-based cancer therapies.
@@ 3.2 Raw data acquisition
Since different data modalities provide complementary biological insights, integrating them can significantly enhance feature learning. In protein research, we identified four key data modalities: sequence, structure, protein pocket, and text. Each modality has distinct characteristics and unique biological information. Below, we introduced the key features of each modality and described the methods used to prepare them for ML applications.
@@
Sequence represents the unique order of amino acids in a protein chain, defining its primary structure. It reveals the specific arrangement of residues that can indicate functional regions, including conserved motifs-such as zinc fingers, H-N-H, and CXXC — that frequently contain key metal-binding residues like cysteine, histidine, aspartic acid, and glutamic acid, which coordinate metal ions critical for structural stability and function [40, 41]. Standard file formats include FASTA, GenBank, and other plain-text formats listing amino acids or nucleotide sequences and can be derived from multiple datasets, including the PDB and UniProt, through their websites or application programming interfaces (APIs). It is also possible to extract protein sequences from a PDB structural data file using tools like Biopython [42].
@@
Structure provides a 3D view of a protein and related metal ions, detailing how its atoms are spatially organized into secondary, tertiary, or quaternary structures, and illustrates the folding pattern and conformational details critical for understanding protein-ligand interactions, including those with metal ions. The secondary structure comprises local folded regions formed by interactions among backbone atoms. The secondary structure of a protein refers to local, repeated arrangements of the amino acid chain, such as alpha-helices and beta-sheets, which are stabilized mainly by hydrogen bonds between the backbone atoms. The tertiary structure represents the overall 3D shape of a single polypeptide chain arising from interactions between the side chains (R groups) of the amino acids. It includes hydrogen bonds, ionic bonds, hydrophobic interactions, and disulfide bridges, which determine the protein’s overall functionality. The quaternary structure applies to proteins composed of multiple polypeptide chains, describing the arrangement and interaction of these subunits. Structural data is typically obtained using X-ray crystallography, NMR spectroscopy, and cryo-electron microscopy (cryo-EM). This information is stored in formats like PDB files or mmCIF, which include 3D coordinates and other structural details. In metal-binding prediction, secondary and tertiary structures are most commonly utilized. Q-BioLip [34] is a recent dataset that provides quaternary structure-based protein-ligand interaction information.
@@
Protein pocket is a 3D cavity within the complex architecture of protein structures which serves as crucial microenvironments for molecular interactions. These indentations can accommodate various ligands, including metal ions, thereby mediating a vast array of protein functions. While pocket data can be extracted from protein structure datasets, several specialized datasets provide detailed pocket data, including PocketDB [38], MetalPDB [31], MetalMine [43], and CavitySpace [39].
@@
Text encompasses the detailed, human-readable information associated with biological entities, such as proteins, found in datasets and scientific literature. Unlike sequences or structures, text provides rich contextual insights, including functional annotations, experimental findings, and disease associations, which complement other data modalities [44]. This modality captures the precise language used to describe protein characteristics, such as metal-binding sites, post-translational modifications, and cellular localization. Resources like UniProt [23], PubMed [45], and PDB present this information in various forms, including free-text entries, abstract summaries, and detailed feature descriptions. Text data is a valuable source for knowledge extraction through natural language processing (NLP) techniques.

### v2

@@ Abstract
In cancer therapeutics, protein-metal binding mechanisms critically govern the pharmacokinetics and targeting efficacy of drugs, thereby fundamentally shaping the rational design of anticancer metallodrugs. While conventional laboratory methods used to study such mechanisms are often costly, low throughput, and limited in capturing dynamic biological processes, machine learning (ML) has emerged as a promising alternative. Despite increasing efforts to develop protein-metal binding datasets and ML algorithms, the application of ML in tumor protein-metal binding remains limited. Key challenges include a shortage of high-quality, tumor-specific datasets, insufficient consideration of multiple data modalities, and the complexity of interpreting results due to the “black box” nature of complex ML models. This paper summarizes recent progress and ongoing challenges in using ML to predict tumor protein-metal binding, focusing on data, modeling, and interpretability. We present multimodal protein-metal binding datasets and outline strategies for acquiring, curating, and preprocessing them for training ML models. Moreover, we explore the complementary value provided by different data modalities and examine methods for their integration. We also review approaches for improving model interpretability to support more trustworthy decisions in cancer research. Finally, we offer our perspective on research opportunities and propose strategies to address the scarcity of tumor protein data and the limited number of predictive models for tumor protein-metal binding. We also highlight two promising directions for effective metal-based drug design: integrating protein-protein interaction data to provide structural insights into metal-binding events and predicting structural changes in tumor proteins after metal binding.
@@
  • Cross-referencing datasets can help address the scarcity of tumor protein-metal-binding data.
@@ 1 Introduction
The design and characterization of metal-based compounds mainly rely on conventional wet-lab methods, such as mass spectrometry, X-ray crystallography, and nuclear magnetic resonance [5, 6, 7, 8], as shown in Fig. 1A. These methods provide valuable insights into the properties of metal-based compounds and the structural basis of their interactions with tumor proteins. However, they are often expensive, time-consuming, and low throughput, and they struggle to capture dynamic biological processes. The emergence of machine learning (ML) presents new opportunities, as ML techniques can efficiently analyze large amounts of complex biological data and build predictive models for protein-metal binding, thereby accelerating the discovery of metal-based drugs. Figure 1D shows three key ML application tasks in tumor protein-metal binding: tumor metalloprotein identification, metal type classification, and binding site prediction. Despite this potential, the application of ML remains limited due to challenges including the scarcity of high-quality datasets, insufficient integration of multimodal information, and a lack of interpretability in existing models.
@@
Figure 1: Interpretable multimodal learning for tumor protein-metal binding. A) The process of transforming bioinformatics data to ML-oriented datasets. B) The data modalities, their combinations, and the commonly-used multimodal learning methods. C) The key roles of model interpretability. D) The primary applications for cancer research. Created with BioRender.
@@
To summarize recent advances, highlight key challenges, and explore promising directions, this paper is organized around three key areas: data resources, multimodal learning approaches, and interpretability techniques. First, we review existing data resources and observe that, while high-quality general protein-metal binding datasets are abundant, those specific to tumor proteins are scarce. Since tumor proteins are not structurally different from general proteins, the learning mechanisms in ML models remain applicable across both types. Therefore, we collect datasets on both general protein-metal binding and tumor proteins. We then propose a strategy to construct tumor-specific datasets by integrating these complementary sources. For example, although the Protein Data Bank (PDB) [9] does not explicitly annotate tumor protein-metal binding interactions, it includes cancer-associated proteins with experimentally resolved metal-binding structures. Identifying and using these entries to extract key biological features can help mitigate current data scarcity.
@@
Next, different data modalities can provide complementary biological insights to enhance feature learning when effectively integrated. To capture this diversity, we identify four key data modalities relevant to tumor protein: sequence, structure, protein pocket, and text, as shown in Fig. 1B. For each modality, we introduce its unique characteristics and present an ML-oriented data processing workflow, including data acquisition, data preprocessing, and feature extraction. This workflow can support researchers in locating, preparing, and adopting raw data for ML model development.
@@
Integrating these heterogeneous modalities into a model is nontrivial and can significantly influence model performance. To better understand the current landscape, we examine existing ML models developed for tumor protein-metal binding. To the best of our knowledge, only three such models have been developed to date - multichannel convolutional neural network (MCCNN) [10], MetalPrognosis [11], and MetalTrans [12] - but they all have limited success in multimodal integration. Therefore, we study multimodal ML methods from broader bioinformatics applications, including 14 recent protein–metal binding models and four protein foundation models, and propose strategies for adapting them to tumor-specific tasks. Building on these insights, we highlight the important value of multimodality for tumor protein-metal binding, suggest promising combinations of data modalities, and provide an overview of general multimodal learning methods, as shown in Fig. 1B.
@@
Finally, in medical and biological applications, predictive accuracy alone is insufficient. ML models should also be interpretable and trustworthy, supporting scientific and clinical understanding. However, the “black box” nature of popular complex ML models limits transparency into their internal decision-making processes. In practice, interpretability is often limited to indirect evaluations, such as performance comparisons with benchmark models, without clearly explaining how specific inputs contribute to predictions. This lack of transparency poses challenges for researchers, particularly those from clinical and biological backgrounds, in assessing model reliability and extracting clinically or biologically relevant insights. To address this issue, we examine interpretability in the context of tumor protein-metal binding from two perspectives: designing models with inherent interpretability and applying post hoc analysis techniques.
@@
Building on these findings, we summarize key challenges and discuss perspectives on potential solutions. Moreover, we identify two emerging directions that can advance the prediction of tumor protein–metal binding. The first involves integrating protein-protein interaction data to provide a structural context for metal-binding events. The second focuses on predicting the structural changes in tumor proteins after metal binding, which can inform the design of more effective metal-based therapies. To our knowledge, this is the first comprehensive study of interpretable multimodal ML for tumor protein-metal binding. Our work synthesizes research advances, identifies ongoing challenges, and offers a forward-looking perspective into future research opportunities. By integrating knowledge from protein structure datasets, multimodal ML, and interpretability analysis, we aim to support the development of biologically grounded, accurate, and clinically applicable ML models for tumor protein-metal binding prediction.
The main contributions of this paper are threefold:
@@
  • We present potential data resources for tumor protein-metal binding, a systematic data processing workflow, and a strategy to construct tumor-specific datasets by integrating protein-metal binding and cancer-related datasets.
@@
  • We identify key protein data modalities, explore their combinations, and review multimodal learning methods focusing on their applicability to tumor protein-metal binding. We also emphasize the importance of interpretability and summarize current approaches that support biological and clinical insight.
@@
  • We offer a forward-looking perspective by proposing two promising research directions, i.e., exploring protein-protein interaction and predicting tumor protein structural changes upon metal binding to inform drug design and discovery.
@@
The remainder of this paper is organized as follows. Section 2 introduces the role of metal ions in tumor proteins, and Section 3 focuses on tumor protein-metal binding data sources and data processing workflow. Section 4 studies related multimodal learning resources and methods, and Section 5 investigates the importance of model interpretability and related methods for interoperability improvement. Section 6 summarizes key challenges and perspectives, highlighting two further directions for advancing tumor protein-metal binding research. Finally, Section 7 concludes this work.
@@ 2 Biological significance of tumor protein-metal binding
These protein-metal interactions form a highly coordinated regulatory system that influences tumor progression, immune response, and cellular metabolism. The biological significance underscores the potential for developing targeted cancer therapies, motivating this work.
@@
3 Tumor protein-metal binding data
Given the limited availability of tumor-specific protein-metal binding datasets, applying ML in this domain remains challenging. This section reviewed existing data resources, identified key data modalities, and proposed strategies to mitigate data scarcity. We introduced publicly available cancer-related and protein-metal binding datasets and discussed how integrating these resources can support the construction of tumor-specific datasets. Additionally, we examine four key data modalities in protein research and present a data processing workflow to prepare these modalities for ML modeling.
3.1 Tumor protein and metal-binding datasets
Several publicly available datasets offer valuable information on cancer biology and protein-metal binding, enabling researchers to investigate the structural and functional roles of metal ions in both tumor and general proteins. As summarized in Table 1 and 2, cancer-related datasets mainly focus on genomic, proteomic, and clinical data relevant to tumor biology, while protein-metal binding datasets contain detailed structural and functional annotations for proteins that coordinate metal ions. Because tumor proteins are not structurally distinct from general proteins, we propose a strategy to integrate tumor protein annotations with metal-binding data, enabling researchers to cross-reference information across specialized datasets. By leveraging multiple data sources, researchers can identify proteins of interest and explore their metal coordination properties and relevance in the context of tumor biology. For example, a protein can be identified through The Cancer Genome Atlas Program (TCGA) [22], and its metal-binding features, such as coordination details, binding motifs, and potential roles, can be explored using UniProt [29]. A case in point is the copper transport protein ATOX1, which exhibits altered RNA transcript levels in various cancers [39]. Researchers can find ATOX1 in TCGA and use its UniProt accession number (O00244) to access its metal-binding information. ATOX1 binds to copper and functions as a chaperone, delivering copper to targets such as P-type ATPases. The dataset also links to studies on ATOX1’s roles in breast and liver cancer, while structural data is accessible via external resources, such as the PDB.
This strategy highlights how integrating datasets such as TCGA and PDB enables researchers to investigate metal-binding proteins, even when explicit metal annotations are missing in tumor datasets, by leveraging cross-referenced structural and functional insights. Looking ahead, we envision a future where an integrated dataset combines gene expression, metal binding, and structural data, streamlining protein-level analysis in cancer. Such a resource would allow researchers to readily query proteins like ATOX1 to assess tumor-specific expression, metal coordination, and structural context, ultimately supporting the discovery of new targets for metal-based cancer therapies.
@@ 3.2 Raw data acquisition
Since different data modalities provide complementary biological insights, integrating them can significantly enhance feature learning and understanding. In protein research, we identify four key data modalities (Fig. 2A): sequence, structure, protein pocket, and text. Each modality has distinct characteristics and unique biological information. Below, we introduce the key features of each modality and describe the methods used to prepare them for ML applications.
@@
Sequence represents the unique order of amino acids in a protein chain, defining its primary structure. It reveals the specific arrangement of residues that can indicate functional regions, including conserved motifs-such as zinc fingers, H-N-H, and CXXC — that frequently contain key metal-binding residues like cysteine, histidine, aspartic acid, and glutamic acid, which coordinate metal ions critical for structural stability and function [40, 41]. Standard file formats include FASTA, GenBank, and other plain text formats that list amino acids or nucleotide sequences. These formats can be derived from multiple datasets, such as the PDB and UniProt, through their websites or application programming interfaces (APIs). It is also possible to extract protein sequences from a PDB structural data file using tools like Biopython [42].
@@
Structure provides a 3D view of a protein and related metal ions, detailing how its atoms are spatially organized into secondary, tertiary, or quaternary structures, and illustrates the folding pattern and conformational details critical for understanding protein-ligand interactions, including those with metal ions. Secondary structures comprise local folded regions formed by interactions among backbone atoms, so secondary structures of a protein refer to local, repeated arrangements of the amino acid chain, such as alpha-helices and beta-sheets, which are stabilized mainly by hydrogen bonds between the backbone atoms. Tertiary structures represent the overall 3D shape of a single polypeptide chain arising from interactions between the side chains (R groups) of the amino acids, including hydrogen bonds, ionic bonds, hydrophobic interactions, and disulfide bridges, which determine the protein’s overall functionality. Quaternary structures apply to proteins composed of multiple polypeptide chains, describing the arrangement and interaction of these subunits. Structural data is typically obtained using X-ray crystallography, NMR spectroscopy, and cryo-electron microscopy (cryo-EM). This information is stored in formats like PDB files or mmCIF, which include 3D coordinates and other structural details. In metal-binding prediction, secondary and tertiary structures are most commonly utilized. Q-BioLip [33] is a recent dataset that provides quaternary structure-based protein-ligand interaction information.
@@
Protein pocket is a 3D cavity within the complex architecture of protein structures that serves as a crucial microenvironment for molecular interactions. These indentations can accommodate various ligands, including metal ions, thereby mediating a vast array of protein functions. While pocket data can be extracted from protein structure datasets, several specialized datasets provide detailed pocket data, including PocketDB [37], MetalPDB [30], MetalMine [43], and CavitySpace [38].
@@
Text encompasses the detailed, human-readable information associated with biological entities, such as proteins, found in datasets and scientific literature. Unlike sequences or structures, text provides rich contextual insights, including functional annotations, experimental findings, and disease associations, which complement other data modalities [44]. This modality captures the precise language used to describe protein characteristics, such as metal-binding sites, post-translational modifications, and cellular localization. Resources like UniProt [29], PubMed [45], and PDB present this information in various forms, including free text entries, abstract summaries, and detailed feature descriptions. Text data is a valuable source for knowledge extraction through natural language processing (NLP) techniques.